# Fluctuations of the free energy of the spherical Sherrington–Kirkpatrick model with heavy-tailed interaction

### v1

@@ 5 Low temperature regime
Then we have
$\int_{\Gamma_{3}}\mathrm{d}y\geq\operatorname{\mathrm{Im}}z_{3}-\operatorname{\mathrm{Im}}\gamma\geq C_{3}N^{-3}.$
Let
$g(z)=\begin{cases}\frac{N}{2b_{N}}[G(z)-G(\gamma)],&|z|>N^{-2}\\ \frac{N}{2b_{N}}[G(z_{2})-G(\gamma)],&|z|<N^{-2}.\end{cases}$
Then $g(z)$ decreases along the curve $\Gamma^{+}$
@@
Since
$|G(\gamma+t)-G(\gamma)|=|\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)t^{k}|=|\sum_{k\geq 2}\frac{1}{k!}G^{(k)}(\gamma)t^{k}|=\Theta(\frac{N}{b_{N}}t^{2})$
for $|t|<N^{-2}$ and $G(\gamma+t)<G(\gamma)$ along the curve,
$\int_{\Gamma_{1}}\exp(\frac{N}{2b_{N}}[G(z_{3})-G(\gamma)])-\exp(\frac{N}{2b_{N}}[G(z_{2})-G(\gamma)])\mathrm{d}y$
$\geq\int_{\Gamma_{1}}\exp(-C_{1}\frac{N}{2b_{N}}N^{-4})-\exp(-C_{2}\frac{N}{b_{N}}N^{-6})\mathrm{d}y\geq C_{3}N^{-3}\frac{N}{2b_{N}}N^{-4}.$
Therefore, this proves that $\frac{1}{N}\log\int_{-\infty}^{\infty}e^{\frac{N}{2b_{N}}(G(\gamma+ti)-G(\gamma))}\to 0$.
@@
where $X$ is the limiting distribution of $\frac{\lambda_{1}}{b_{N}}.$ We conclude the proof of the desired statement. ∎
@@ Appendix A Proofs of technical lemmas
for $k\leq M$ and $\gamma=\frac{1}{4M+2}$. Moreover, applying this for every row and we have
@@
This implies the first case. Case 2: $0<\alpha\leq 1$.
@@
$\mathbb{P}(Cb_{N}-b_{N}N^{-\epsilon}<\lambda_{1}<Cb_{N})$
$\leq\mathbb{P}(Cb_{N}-b_{N}N^{-\epsilon}<M_{0}(1+O(b_{N}^{-1/8}))<Cb_{N})+\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))$
@@
for some $\epsilon^{\prime}>0$.
@@
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>N^{4/\alpha+\epsilon})\leq O(N^{-\alpha\epsilon/4}).$
@@
Since
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>b_{N}^{2}N^{\epsilon})\leq O(N^{-\alpha\epsilon/8}),$
with high probability,
@@
Thus, with high probability,
@@
for some non-dengerate random variable Y.
@@
This implies
@@ Appendix B Statistical property for T
[…]
@@
for any cases. Now let us get back to calculate the limit of $\mathbb{E}(T^{2})$. To calculate this, we calculate
@@
[…]

### v2

@@ 5 Low temperature regime
In Lemma 5.3, we will show that for any real valued function $f(z)$ on $\Gamma^{+}$ that decreases along the curve as $z$ moves from $\gamma$ to $\infty$- ∞, $\int_{\Gamma^{+}}e^{f(z)}\mathrm{d}y\geq 0$. With this lemma, if we let
$g(z)=\begin{cases}\frac{N}{2b_{N}}[G(z)-G(\gamma)],&|z|>N^{-2},\\ \frac{N}{2b_{N}}[G(z_{2})-G(\gamma)],&|z|<N^{-2},\end{cases}$
Then $g(z)$ decreases along the curve $\Gamma^{+}$ and
@@
We need to find a lower bound for the right-side of the equation above. Considering again the Taylor expansion of $G(\gamma+z)$ with the estimates in Lemma 5.2, we get
$G(z_{2})-G(\gamma)\leq-C_{1}\frac{N^{1-2\epsilon}}{b_{N}}N^{-4},\quad G(z_{3})-G(\gamma)\geq-C_{2}\frac{N}{b_{N}}N^{-6}$
for some constants $C_{1},C_{2}>0$. Since
$\int_{\Gamma_{1}}\mathrm{d}y\geq\operatorname{\mathrm{Im}}z_{3}-\operatorname{\mathrm{Im}}\gamma\geq C_{3}N^{-3}$
for some constant $C_{3}>0$, we find that
$\int_{\Gamma_{1}}\exp(\frac{N}{2b_{N}}[G(z_{3})-G(\gamma)])-\exp(\frac{N}{2b_{N}}[G(z_{2})-G(\gamma)])\mathrm{d}y\geq C_{4}N^{-3}\frac{N}{2b_{N}}\frac{N^{1-2\epsilon}}{b_{N}}N^{-4}$
for some constant $C_{4}>0$. This proves the desired lower bound for $K$ and hence the desired claim.
@@
where $X$ is the limiting distribution of $\frac{\lambda_{1}}{b_{N}}$. Finally, $\lim_{N\to\infty}\mathbb{P}(F_{2})$ can be easily checked from Proposition 3.2. This concludes the proof of Proposition 5.1. ∎
@@ Appendix A Proofs of technical lemmas
for $k\leq M$ and $\gamma=\frac{1}{4M+2}$. We apply this for every row and we have
@@
This implies the first case.
Case 2: $0<\alpha\leq 1$.
@@
$\mathbb{P}(Cb_{N}(1-N^{-\epsilon})<\lambda_{1}<Cb_{N}(1+N^{-\epsilon}))$
$\leq\mathbb{P}(Cb_{N}(1-N^{-\epsilon})<M_{0}(1+O(b_{N}^{-1/8}))<Cb_{N}(1+N^{-\epsilon}))+\mathbb{P}(\lambda_{1}\neq M_{0}(1+O(b_{N}^{-1/8})))$
@@
for some $\epsilon^{\prime}>0$. The special case $C=\frac{1}{2\beta}$ is our desired result.
@@
Since above lemma and $b_{N}<N^{2/\alpha+\epsilon/10}$ with high probability due to Lemma 3.12,
$\mathbb{P}(\sum_{i}\lambda_{i}^{2}>b_{N}^{2}N^{\epsilon})=O(N^{-\alpha\epsilon/8})$
holds with high probability.
Hence, with high probability, we obtain
@@
and this implies
@@
for some non-degenerate random variable Y.
@@
Using the same method of proving Lemma 4.3, this implies
@@ Appendix B Statistical property for T
$\mathbb{E}(|Y|)=\int_{0}^{b_{N}/2\beta}\mathbb{P}(|Y|>u)\mathrm{d}u=\begin{cases}\frac{(2\beta)^{\alpha}}{1-\alpha}\frac{b_{N}}{2\beta}\frac{2}{N(N+1)}(1+o(1))&\alpha<1\\ O(1)&\textit{otherwise}\end{cases}$
@@
for any cases. Now, let us get back to the calculation of the limit of $\mathbb{E}(T^{2})$. To calculate this, we calculate
@@
$\lim_{N\to\infty}\mathbb{E}((\frac{b_{N}}{2\beta})^{-k-l}(\sum\lambda_{i}^{k})(\sum\lambda_{i}^{l}))=\begin{cases}\frac{(2(2\beta)^{\alpha})^{2}kl}{(k-\alpha)(l-\alpha)}+\frac{2(2\beta)^{\alpha}(k+l)}{k+l-\alpha}&k,l\textit{ are even}\\ 0&\textit{otherwise}\end{cases}$
@@ Acknowledgement
Conflict of interest
The authors have no relevant financial or non-financial interests to disclose.
Data Availability
No datasets were generated or analysed during the current study.